A roulette wheel is constructed in such a way that the numbers from 1 to 10 can fall into it and the probability of dropping a number (k + 1) is 2 times greater than dropping the number k (for k = 1; 2; 3; ... 9). Find the probability that: a falls the number 1; b to drop the number 10. Give the ending as a decimal number rounded to 3 digits after the decimal point.


1
Expert's answer
2021-03-26T08:36:14-0400

"\text{Let }X\text{ be our random variable (the number that falls)}.\\
x\text{ --- probability that the number 1 falls}\\
2x\text{ --- probability that the number 2 falls}\\
4x\text{ --- probability that the number 3 falls}\\
x+2x+4x+\ldots=1\\
\text{The left side is the sum of the geometric progression with }n=10,\ q=2,\ b_1=1.\\
S_{n}=\frac{b_1(1-q^n)}{1-q}=\frac{1(1-2^{10})}{1-2}=1023.\\
1023x=1\\
x=\frac{1}{1023}\approx 0.001."

a. So the probability that the number 1 falls is "0.001".



The probability that the number 2 falls is "\frac{2}{1023}\approx 0.002."

"b_{10}=b_1q^9=1\cdot 2^9=512."

b. The probability that the number 10 falls is "\frac{512}{1023}\approx 0.500."